Please solve using minitab and explain steps. Please show all work
9.14 The quality-control manager at a light emitting diode (LED) factory needs to determine whether the mean life of a large shipment of LEDs is equal to 50,000 hours. SELF TEST
mple Tests The population standard deviation is 1,500 hours, A of 64 LEDs indicates a sample mean life of 49.875 a. At the 0.05 level of significance, is there evidence that ours. the life is different from 50,000 hours?
Answer #1

Claim: Mean life of shipment of LED's is 50000 hours that is μ = 50000

Step 1: To write the null and alternative hypothesis

Ho : μ = 50000(claim) and H1 : μ 50000

Step 2: Test statistics

Here the population standard deviation is known so the z test is applicable.

The formula of test statistics is

Z = rac{ar x - mu }{sigma /sqrt{n}}

Where r-sample mean = 49875. σ-population standard deviation 1500

μ = 50000 . n-sample size = 64

μ-49875-50000 σ/νη Z = x ー-0.67 1500/V64

Step 3: Critical value

Here the alternative hypothesis contains not equal to sign so the test is two tailed test.

There are 2 critical values both are same just opposite in sign.

Here alpha (level of significance) = 0.05

Alpha/2 = 0.05/2 = 0.025

Search 0.025 in the middle body of table and then take the corresponding z score. The corresponding z score is -1.96

So the Critical values are -1.96 and 1.96

Step 4: Decision rule

If test statistics falls between both the critical values then we fail to reject the null hypothesis otherwise we reject the null hypothesis.

Here test statistics -0.67 falls between -1.96 and 1.96, so we fail to reject the null hypothesis.

Step 5: Conclusion

Fail to reject the null hypothesis, that is there is no sufficient evidence that the mean life is different from 50000 hours.